# Probing Binary Lens Caustics with Gravitational Waves: A Uniform Approximation Approach

### v2

@@ I Introduction
One particularly fascinating aspect of gravitational lensing is the substantial magnification observed near caustics. This effect has been thoroughly explored in the context of EM waves. Due to the extraordinarily high amplification—often exceeding $10^{4}$ times—objects such as individual stars and distant galaxies can be detected at redshifts far beyond the reach of current detector sensitivity for unlensed observations [11]. This exceptional capability has opened up new possibilities for observing faint and distant astrophysical phenomena, allowing constraints on source size and structure while providing insigths into the accretion disks and broad-line regions of quasars [12, 13, 14].
Recent advancements in GW interferometry have brought these detectors to the level of sensitivity required to observe gravitational lensing effects in GWs. During the first three observing runs, the global network of GW detectors—Advanced LIGO [15], Advanced Virgo [16], and KAGRA [17]—successfully detected 90 compact binary coalescence events [18, 19, 20, 21]. A range of phenomena, from cosmological to astrophysical scales, could become accessible through the lensing of GWs. Similarly, it is expected that gravitational lensing of GWs could help identify non-luminous objects that bend the path of the waves, producing effects like those seen with light.
@@
When a GW source approaches a caustic, the signal detected by the interferometer is expected to show significant amplification. This phenomenon is well-studied in the field of EM waves but remains largely unexplored in the context of GWs. It is intriguing from an astrophysical perspective and is currently a subject of significant interest in ongoing research [24, 25]. However, near the caustic, the geometric optics (GO) approximation no longer provides an accurate description [26, 27, 28, 29]. Specifically, as the source approaches a caustic point, the time delay between the merging (or coalescing) images becomes vanishingly small. This implies that achieving the GO limit would require infinitely high GW frequencies. Full-wave modeling provides greater accuracy than the GO approximation. However, it becomes computationally intensive and demands significant computational resources, particularly for complex mass distributions. These limitations underscore the need for alternative strategies. Asymptotic methods, which bridge the gap between GO and full-wave approaches, offer a promising solution by capturing the essential physics without requiring exhaustive computations.
@@
One crucial difference between EM waves and GWs lies in the scale of their wavelengths. For EM waves, the wavelength is typically much smaller, so the region near the caustic where the GO approximation breaks down is very small, often negligible, especially when considering galaxy-scale masses. In this regime, where wavelengths are very short and masses are high, the geometric optics approximation remains sufficiently accurate for analyzing the wave field down to very close distances from the caustic. For GWs, however, the wavelength is much larger, typically $\gtrsim 10^{6}$ meters in the LIGO-Virgo-KAGRA (LVK) frequency band [30]. As a result, the region near the caustic where the geometric optics approximation breaks down is significantly broader and may be substantial compared to the characteristic dimension of the lens. In this context, the GO approximation fails to provide an accurate description, and a wave-optics approach is required to capture the behavior of the waves near the caustic.
@@
As an alternative to the GO approximation, one of the most notable methods is the Transitional Approximation [28], which describes the wave field near a caustic in terms of the Airy function [31]. This solution provides a continuous representation across the fold, avoiding the divergences at the caustic seen in GO. The key is that it expands the phase to the cubic term, whereas the divergence in GO arises from the quadratic term, where the second derivative vanishes. The approximation using the Airy function is widely employed in the literature [32, 24, 25]. However, while it provides an excellent local solution in the immediate vicinity of the singular region near the fold singularity, it is not valid at larger distances and does not align with the GO approximation far from the fold, where the latter is certainly valid. The two solutions converge only in the limit of infinitely small wavelengths (a good approximation for EM waves) or exceptionally high lens masses, on the order of a galaxy’s mass [25]. To fully characterize this phenomenon and provide a comprehensive description valid across a wide range of parameters, we introduce the caustic width, $d_{c}$, as a physical parameter that enables the study and analysis of various limiting cases.
@@
In this paper, we introduce the Uniform Approximation, an alternative asymptotic method that accurately models the wave field near the caustic fold while also smoothly connecting with the GO solution far from the caustic. This approach offers a consistent and versatile solution, particularly suited for integration into GW lensing pipelines. Importantly, it relies on analytical functions, specifically the Airy function and its derivative, and avoids extensive computations. This makes it particularly advantageous for scenarios where the source crosses the fold caustic.
@@ II Microlensing on BML
where $M_{L}$ is the mass of the lens, $d_{LS}$ is the angular diameter distance between the source and the lens, and $d_{L}$ and $d_{S}$ are the angular diameter distances to the lens and source at redshifts $z_{L}$ and $z_{S}$, respectively. $R_{\rm E}$ represents the characteristic length scale on the lens plane, is proportional to $\sqrt{M_{L}}$ and is typically much smaller than the cosmological distances $d_{LS}$, $d_{L}$, and $d_{S}$. This enables the lens mass to be projected onto a lens plane. In the thin lens approximation, gravitational waves propagate freely outside the lens, interacting only with a two-dimensional gravitational potential at the lens plane, where the lensing effect is ultimately captured in the transmission factor $F$ [32]. Another important length scale is the distance between the binary masses, which will be introduced below.
@@ II.1 Wave approach to gravitational lensing
Next, we will consider a binary mass system as the lens. We can consider both masses as point-like when their physical sizes are much smaller than the Einstein radius of each, as in the case of binary black holes, stars, dense dark matter clumps, and similar compact objects. The single point mass lens (PML) model has been widely used in the literature to interpret both electromagnetic [36, 37, 32, 38] and GW lensing [39, 26, 33, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. However, the binary mass lens (BML) model has been less commonly applied in comparison [58, 59].
@@
where ${\cal R_{\rm S}}=2GM_{Lz}/c^{2}$ is the Schwarzschild radius, and $M_{Lz}=M_{L}(1+z_{L})$ represents the redshifted mass of the lens. The diffraction integral (3), expressed in terms of dimensionless quantities, is given by:
$F(w,\mathbf{y})=\frac{w}{2\pi\mathrm{i}}\int e^{\mathrm{i}wT(\mathbf{x},\mathbf{y})}\,\mathrm{d}^{2}\mathbf{x}$ (8)
@@ III Transitional Asymptotics
The most intriguing scenario arises when the source is near the caustic, as it produces the highest signal amplifications, occurring precisely where multiple rays converge. We aim to describe the wave properties near caustics, where the GO approximation fails, using asymptotic methods rather than relying on the full-wave approach. Near a fold, two stationary points become closely spaced, making it impossible to separate their contributions using the simplest form of the stationary phase method, as applied in the GO limit. In the BML model under consideration, one of these points corresponds to a saddle point within the interior region, while the other represents a minimum outside it (see Fig. 2). These points are denoted as $S$ and $M$, respectively, with the condition that the time delays satisfy $T_{M}<T_{S}$. As these points merge, the phase extremum in the Kirchhoff integral becomes degenerate, being stationary to a higher order: both the linear and quadratic terms vanish. See the evolution of the time delay in Fig. 6 in Appendix A. The solution is to approximate the phase by a polynomial expanded up to a nonzero cubic term, preserving the stationary-point structure, and then evaluate the resulting integral, which is more complex than a Gaussian, using special functions.
@@
$\mathrm{Ai}(u)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\exp{[\mathrm{i}\left(u\tilde{t}+\tilde{t}^{3}/3\right)]}\,\mathrm{d}\tilde{t}.$
provides an accurate description of wave behavior near the fold catastrophe. For $u<0$ (inside the caustic), it exhibits oscillations due to the interference of the two merging stationary points, while for $u>0$ (outside the caustic), it decays exponentially. At $u=0$, it smoothly transitions between these regimes, offering a Transitional Approximation (TA) across the fold [28] (see e.g. Refs. [32, 63, 64, 26] where the Airy function has been applied to lensing at the fold). Finally, the transmission factor near a fold in the TA framework is given by
@@
where $C=(2\pi/|T_{11}|)^{1/2}$ and $\rho_{c}=|T_{222}|/2$ are expressed through derivatives evaluated at the fold, and hence depend parametrically on the parameter $b$. For the symmetric configuration we obtain, $C=\sqrt{\pi}$ and
@@
It is important to note that Eq. (16) is specifically derived for the fold catastrophe and cannot be directly applied to other caustic types. However, the methodology used to obtain this result can be extended to higher-order catastrophes, such as the cusp, by mapping the time delay to their corresponding standard catastrophe functions, which serve as the dominant contributions to the diffraction integral close to caustics. In the present work, we will not pursue this generalization.
@@ IV Uniform Asymptotics
The wave field obtained within the TA framework, expressed through the Airy function, remains valid only in the immediate vicinity of the fold singularity. Further away, the time delay function deviates from the cubic polynomial associated with the fold catastrophe (14), leading to the breakdown of the TA approximation. In the short-wavelength limit, the wave field can instead be described using geometric rays far from the caustics. Yet, the GO and TA solutions do not smoothly connect, as they rely on Taylor expansions around different stationary points. In the GO approximation, these points are determined by quadratic polynomials, leading to a diverging field at the fold. In contrast, the TA framework is based on a cubic polynomial with distinct extrema of the time delay function, yielding a finite solution at the fold. A more robust approach would be to construct an asymptotic solution that remains valid across the singularity while also being expressed in terms of the geometric ray parameters that describe the wave field far from it. This conceptual limitation—where the TA does not inherently involve variables defined by the physical rays—can be overcome using the mathematical framework of Uniform Approximation (UA), initially introduced by Chester et al. [65] and later applied to wave theory by Kravtsov [66] and Ludwig [67].
@@
which is essentially the Airy function but with the explicit large parameter $\kappa$. The ”caustic index” 2/3 and the ”focusing index” 1/6 are associated with simple caustics of the fold type [62]. With the parametrization $w=\kappa/\rho_{c}$, $\tilde{y}_{2}=\rho_{c}\,u$, the TA solution from Eq. (16) reduces to the fold function (18) with a prefactor determined by $b$:
@@
At the fold, the arguments of $\mathrm{Ai}$ and $\mathrm{Ai}^{\prime}$ are zero, as $\tau$ vanishes, while the GO amplitudes $\mu_{j}$ diverge. However, the prefactor of $\mathrm{Ai}$ involves an indeterminate form of type $0\times\infty$ that converges to a finite limit, and similarly, the prefactor of $\mathrm{Ai}^{\prime}$ follows an ×$\infty\times 0$ form that also remains finite. As a result, the singularities cancel out.
Outside the caustic region (often referred to as the caustic shadow) for $u>0$, the two GO rays become complex, resulting in time delays that form a complex conjugate pair: $T_{\pm}=T_{r}\pm\mathrm{i}T_{i}$. To ensure a physically meaningful solution, we discard the ray with a negative imaginary part, $T_{i}<0$, that grows as we move further from the caustic. The second ray, which has a positive imaginary part, $T_{i}>0$, leads to an asymptotically decaying solution away from the caustic and should match the asymptotic behavior of the fold function for positive arguments:
@@
$F_{\rm UA}=2\sqrt{\pi}\,|\mu_{+}|^{1/2}(w\tau_{+})^{1/6}\times\\ \rm{Ai}\left[(w\tau_{+})^{2/3}\right]e^{\mathrm{i}w\tau_{0}^{+}-\mathrm{i}\phi^{+}},$ (29)
where $\tau_{+}=(3/2)\,\mathrm{Im}(T_{+})$, $\tau_{0}^{+}=\mathrm{Re}(T_{+})$, and $\phi^{+}=\rm{Arg}(\det\hat{K}_{+})/2$ represents the argument of the complex determinant of the Hessian matrix defined in Eq. (12).
@@ V.1 Caustic width
Twice this distance, $2d_{c}$, can be referred to as the ”caustic width” (see Fig. 3). Recalling that we previously defined $w=2\pi ft_{M}=4\pi{\cal R}_{\rm S}/\lambda$, we can conclude that the caustic width scales with the wavelength as $\sim\lambda^{2/3}$, and with the lens mass as $\sim M_{Lz}^{-2/3}$. Based on this, we suggest the following expression for the caustic width (normalized to the Einstein radius):
$2d_{c}$ $\simeq$ $0.655\,\rho_{c}^{1/3}\left(\frac{\lambda}{2{\cal R}_{\rm S}}\right)^{2/3}$ (31)
@@
We can now highlight several important conclusions: (i) On the one hand, the introduction of the caustic width allows us to estimate a new characteristic length scale $d_{c}$ associated with wave-optics effects. Due to the universality of the time-delay expansion near a fold, its dependence on the wavelength $\sim\lambda^{2/3}$ and the lens mass $\sim M_{Lz}^{-2/3}$ is also universal. Only the numerical coefficient in front is specific to the particular fold and the mass distribution of the lens. (ii) On the other hand, as we will see in the next chapter, the maximum amplification can be expressed in terms of the caustic width. (iii) Finally, $d_{c}$ defines the region around the fold where the two overlapping rays can no longer be distinguished due to diffraction. That is, the rays can be physically distinguished only when $|\tilde{y}_{2}|>d_{c}$ within the interior zone.

### v3

@@ I Introduction
One particularly fascinating aspect of gravitational lensing is the substantial magnification observed near caustics. This effect has been thoroughly explored in the context of EM waves. Due to the extraordinarily high magnification—often exceeding $10^{4}$ times—objects such as individual stars and distant galaxies can be detected at redshifts far beyond the reach of current detector sensitivity for unlensed observations [11]. This exceptional capability has opened up new possibilities for observing faint and distant astrophysical phenomena, allowing constraints on source size and structure while providing insights into the accretion disks and broad-line regions of quasars [12, 13, 14].
Recent advancements in GW interferometry have brought these detectors to the level of sensitivity required to observe gravitational lensing effects in GWs. During the first three observing runs, the global network of GW detectors—Advanced LIGO [15], Advanced Virgo [16], and KAGRA [17]—achieved a steadily increasing detection rate of compact binary coalescences, resulting in the observation of 90 events by the end of O3 [18, 19, 20, 21]. This significant increase in detection rate has markedly enhanced the probability of observing gravitational lensing signatures in GW signals [22, 23]. A range of phenomena, from cosmological to astrophysical scales, could become accessible through the lensing of GWs. These include, for instance, the measurement of galaxy velocity dispersions, constraints on dark matter and binary formation channels, the detection of intermediate-mass black holes, and novel tests of general relativity, among others (see, e.g., the references listed in [22, 23]).
@@
When a GW source approaches a caustic, the signal detected by the interferometer is expected to show significant amplification. This phenomenon is well-studied in the field of EM waves but remains largely unexplored in the context of GWs. It is intriguing from an astrophysical perspective and is currently a subject of significant interest in ongoing research [26, 27]. However, near the caustic, the geometrical optics (GO) approximation no longer provides an accurate description [28, 29, 30, 31]. Specifically, as the source approaches a caustic point, the time delay between the merging (or coalescing) images becomes vanishingly small. This implies that achieving the GO limit would require infinitely high GW frequencies. Full-wave modeling provides greater accuracy than the GO approximation. However, it becomes computationally intensive and demands significant computational resources, particularly for complex mass distributions. These limitations underscore the need for alternative strategies. Asymptotic methods, which bridge the gap between GO and full-wave approaches, offer a promising solution by capturing the essential physics without requiring exhaustive computations.
@@
One crucial difference between EM waves and GWs lies in the scale of their wavelengths. For EM waves, the wavelength is typically much smaller, so the region near the caustic where the GO approximation breaks down is very small, often negligible, especially when considering galaxy-scale masses. In this regime, where wavelengths are very short and masses are high, the geometrical optics approximation remains sufficiently accurate for analyzing the wave field down to very close distances from the caustic. For GWs, however, the wavelength is much larger, typically $\gtrsim 10^{6}$ meters in the LIGO-Virgo-KAGRA (LVK) frequency band [32]. As a result, the region near the caustic where the geometrical optics approximation breaks down is significantly broader and may be substantial compared to the characteristic dimension of the lens. In this context, the GO approximation fails to provide an accurate description, and a wave-optics approach is required to capture the behavior of the waves near the caustic.
@@
As an alternative to the GO approximation, one of the most notable methods is the Transitional Approximation [30], which describes the wave field near a caustic in terms of the Airy function [33]. This solution provides a continuous representation across the fold, avoiding the divergences at the caustic seen in GO. The key difference lies in the expansion of the phase of the wave field as a function of the spatial coordinate near the caustic: the standard GO approximation expands the phase only up to the quadratic (second-order) term, which causes divergences when the second derivative vanishes at the caustic. In contrast, the Transitional Approximation includes the cubic (third-order) term in the phase expansion, which regularizes the solution and accurately captures the wave behavior near the fold. The approximation using the Airy function is widely employed in the literature [34, 26, 27]. However, while it provides an excellent local solution in the immediate vicinity of the singular region near the fold singularity, it is not valid at larger distances and does not align with the GO approximation far from the fold, where the latter is certainly valid. The two solutions converge only in the limit of infinitely small wavelengths (a good approximation for EM waves) or high lens masses, comparable to that of a galaxy [27]. To fully characterize this phenomenon and provide a comprehensive description valid across a wide range of parameters, we introduce the caustic width, $d_{c}$, as a physical parameter that enables the study and analysis of various limiting cases.
@@
In this paper, we introduce the Uniform Approximation: an alternative asymptotic method that accurately models the wave field near a fold caustic (also known as a fold catastrophe in the context of catastrophe theory [35]) and smoothly connects with the GO solution far from the caustic. This approach offers a consistent and versatile solution, particularly suited for integration into GW lensing pipelines. Importantly, it relies on analytical functions, specifically the Airy function and its derivative, and avoids extensive computations. This makes it particularly advantageous for scenarios where the source crosses the fold caustic.
@@ II Microlensing on BML
where $M_{L}$ is the mass of the lens, $d_{LS}$ is the angular diameter distance between the source and the lens, and $d_{L}$ and $d_{S}$ are the angular diameter distances to the lens and source at redshifts $z_{L}$ and $z_{S}$, respectively. $R_{\rm E}$ represents the characteristic length scale on the lens plane, is proportional to $\sqrt{M_{L}}$ and is typically much smaller than the cosmological distances $d_{LS}$, $d_{L}$, and $d_{S}$. This enables the lens mass to be projected onto a lens plane. In the thin lens approximation, gravitational waves propagate freely outside the lens, interacting only with a two-dimensional gravitational potential at the lens plane, where the lensing effect is ultimately captured in the transmission factor $F$ [34]. Another important length scale is the distance between the binary masses, which will be introduced below. For a schematic diagram of lensing by a binary mass lens, we refer the reader to Fig. 1.
@@ II.1 Wave approach to gravitational lensing
Next, we will consider a binary mass system as the lens. We can treat each lensing mass as point-like when its physical radius is much smaller than its respective Einstein radius, as in the case of binary black holes, stars, dense dark matter clumps, and similar compact objects. The single point mass lens (PML) model has been widely used in the literature to interpret both electromagnetic [39, 40, 34, 41] and GW lensing [42, 28, 36, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. However, the binary mass lens (BML) model has been less commonly applied in comparison [61, 62].
@@
where $R_{\rm S}=2GM_{Lz}/c^{2}$ is the Schwarzschild radius, and $M_{Lz}=M_{L}(1+z_{L})$ represents the redshifted mass of the lens. The diffraction integral (3), expressed in terms of dimensionless quantities, is given by:
$F(w,\mathbf{y})=\frac{w}{2\pi\mathrm{i}}\iint e^{\mathrm{i}wT(\mathbf{x},\mathbf{y})}\,\mathrm{d}^{2}\mathbf{x}$ (8)
@@
We note that, in our analysis, the lens configuration can be considered static during the wave passage because the orbital period of the lens binaries is much longer than the timescale associated with the motion of the source binaries (i.e., the GW period and chirp duration). This difference in timescales reflects the vastly different length scales involved: the lens binary separation is on the order of $R_{\rm E}$, while that of the source is on the order of $R_{\rm S}$.
@@ III Transitional Asymptotics
The most intriguing scenario arises when the source is near the caustic, as it produces the highest signal amplification, occurring precisely where multiple rays converge. We aim to describe the wave properties near caustics, where the GO approximation fails, using asymptotic methods rather than relying on the full-wave approach. Near a fold, two stationary points become closely spaced, making it impossible to separate their contributions using the simplest form of the stationary phase method, as applied in the GO limit. In the BML model under consideration, one of these points corresponds to a saddle point within the interior region, while the other represents a minimum outside it (see Fig. 2). These points are denoted as $S$ and $M$, respectively, with the condition that the time delays satisfy $T_{M}<T_{S}$. As these points merge, the phase extremum in the Kirchhoff integral becomes degenerate. This means that the first and second derivatives of the phase—or equivalently, the first- and second-order terms in its Taylor expansion around the stationary point—both vanish. The evolution of this time delay is illustrated in Fig. 6 in Appendix A. To accurately model this, we must approximate the phase using a polynomial expanded up to a non-zero third-order term. This higher-order approximation is crucial for preserving the stationary-point structure. The resulting integral, being more complex than a simple Gaussian integral, then requires evaluation using special functions.
@@
$\mathrm{Ai}(u)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\exp{[\mathrm{i}\left(u\tilde{t}+\tilde{t}^{3}/3\right)]}\,\mathrm{d}\tilde{t}$
provides an accurate description of wave behavior near the fold singularity. For $u<0$ (inside the caustic), it exhibits oscillations due to the interference of the two merging stationary points, while for $u>0$ (outside the caustic), it decays exponentially. At $u=0$, it smoothly transitions between these regimes, offering a Transitional Approximation (TA) across the fold [30] (see e.g. Refs. [34, 65, 66, 28] where the Airy function has been applied to lensing at the fold). Finally, the transmission factor near a fold in the TA framework is given by
@@
where $C=(2\pi/|T_{11}|)^{1/2}$ and $\rho_{c}=|T_{222}|/2$ are expressed through derivatives evaluated at the fold, and hence depend parametrically on the parameter $b$. For the symmetric configuration we obtain $C=\sqrt{\pi}$ and
@@
It is important to note that Eq. (16) is specifically derived for the fold catastrophe and cannot be directly applied to other caustic types. However, the methodology used to obtain this result can be extended to higher-order catastrophes, such as the cusp, by mapping the time-delay function to the corresponding canonical diffraction integrals—standard forms in catastrophe theory (Airy for fold, Pearcey for cusp, etc.), which dominate the wavefield near caustic singularities [35]. In the present work, we will not pursue this generalization.
@@ IV Uniform Asymptotics
The wave field obtained within the TA framework, expressed through the Airy function, remains valid only in the immediate vicinity of the fold singularity. Further away, the time delay function deviates from the cubic polynomial associated with the fold catastrophe [Eq. (14)], leading to the breakdown of the TA approximation. In the short-wavelength limit, the wave field can instead be described using geometric rays far from the caustics. Yet, the GO and TA solutions do not smoothly connect, as they rely on Taylor expansions around different stationary points. In the GO approximation, these points are determined by quadratic polynomials, leading to a diverging field at the fold. In contrast, the TA framework is based on a cubic polynomial with distinct extrema of the time delay function, yielding a finite solution at the fold. A more robust approach would be to construct an asymptotic solution that remains valid across the singularity while also being expressed in terms of the geometric ray parameters that describe the wave field far from it. This conceptual limitation—where the TA does not inherently involve variables defined by the physical rays—can be overcome using the mathematical framework of Uniform Approximation (UA), initially introduced by Chester et al. [67] and later applied to wave theory by Kravtsov [68], Ludwig [69, 70], and Berry [71].
@@
which is essentially the Airy function but with the explicit large parameter $\kappa$. The values of the indices 2/3 and 1/6 are associated with simple caustics of the fold type [35]. With the parametrization $w=\kappa/\rho_{c}$, $\tilde{y}_{2}=\rho_{c}\,u$, the TA solution from Eq. (16) reduces to the fold function (18) with a prefactor determined by $b$:
@@
At the fold, the arguments of $\mathrm{Ai}$ and $\mathrm{Ai^{\prime}}$ are zero, as $\tau$ vanishes, while the GO amplitudes $\mu_{j}$ diverge. The prefactors of $\mathrm{Ai}$ and $\mathrm{Ai^{\prime}}$ involve indeterminate forms of type $0\times\infty$ and ×$\infty\times 0$, respectively, but explicit evaluation shows that both expressions remain finite. Thus, the apparent singularities cancel out.
Outside the caustic region (often referred to as the caustic shadow) for $u>0$, the two GO rays become complex, resulting in time delays that form a complex conjugate pair: $T_{\pm}=T_{r}\pm\mathrm{i}T_{i}$. To ensure a physically meaningful solution, we discard the ray with a negative imaginary part, $T_{i}<0$, that grows as we move farther from the caustic. The second ray, which has a positive imaginary part, $T_{i}>0$, leads to an asymptotically decaying solution away from the caustic and should match the asymptotic behavior of the fold function for positive arguments:
@@
$F_{\rm UA}=2\sqrt{\pi}\,|\mu_{+}|^{1/2}(w\tau_{+})^{1/6}\times\\ \mathrm{Ai}[(w\tau_{+})^{2/3}]\,e^{\mathrm{i}w\tau_{0}^{+}-\mathrm{i}\phi^{+}},$ (29)
where $\tau_{+}=(3/2)\,\mathrm{Im}(T_{+})$, $\tau_{0}^{+}=\mathrm{Re}(T_{+})$, and $\phi^{+}=\mathrm{Arg}(\det\hat{K}_{+})/2$ represents the argument of the complex determinant of the Hessian matrix defined in Eq. (12).
@@ V.1 Caustic width
Twice this distance, $2d_{c}$, can be referred to as the ”caustic width” (see Fig. 3). Recalling that we previously defined $w=2\pi ft_{M}=4\pi R_{\rm S}/\lambda$, we can conclude that the caustic width scales with the wavelength as $\sim\lambda^{2/3}$, and with the lens mass as $\sim M_{Lz}^{-2/3}$. Based on this, we suggest the following expression for the caustic width (normalized to the Einstein radius):
$2d_{c}$ $\simeq$ $0.655\,\rho_{c}^{1/3}\left(\frac{\lambda}{2R_{\rm S}}\right)^{2/3}$ (31)
@@
We can now highlight several important conclusions: (i) On the one hand, the introduction of the caustic width allows us to estimate a new characteristic length scale $d_{c}$ associated with wave-optics effects. Due to the universality of the time-delay expansion near a fold, its dependence on the wavelength $\sim\lambda^{2/3}$ and the lens mass $\sim M_{Lz}^{-2/3}$ is also universal. Only the numerical coefficient in front is specific to the particular fold and the mass distribution of the lens. (ii) On the other hand, as we will see in Sec. V.2, the maximum amplification can be expressed in terms of the caustic width. (iii) Finally, $d_{c}$ defines the region around the fold where the two overlapping rays can no longer be distinguished due to diffraction. That is, the rays can be physically distinguished only when $|\tilde{y}_{2}|>d_{c}$ within the interior zone.